### v3

@@ 4.2 Market Environments
The financial data is processed into the standard gym-style market environments to build a standard practical environment. This part describes the market environments with near-real market constraints and LLM-generated signals. To address the sampling bottleneck of the training stage, we develop massively parallel market environments on GPUs.
@@ 4.2.1 Market Environment with Trading Constraints
The operations in the FinRL market environment include:
@@ 4.2.2 Massively Parallel Simulations on GPUs
  • NN is the number of parallel SubEnvs.
@@
Improved Sampling Speed with Massively Parallel Environments on GPU. We evaluated the sampling speed measured in samples per second using vectorized environments for stock trading. We used the PPO agent and the OHLCV data of $30$ constituent stocks in the Dow Jones index, from 2020-01-01 to 2022-01-01. NVIDIA A100 GPU is used. The numbers of parallel environments vary from $1$, 22, 44, $\ldots$, to $2,048$. As shown in Fig. 4, the average sampling speed with $2,048$ parallel environments is $227,212.54$ samples per second. The sampling speed is improved by $1,649.93\times$ compared with a single environment. The sampling speed scales approximately linearly with the number of parallel environments. The results show the effectiveness of massively parallel simulation in improving sampling speed in FinRL Contest tasks.
@@ 5.1.1 Market Index and Mean-Variance Method
  • Mean-variance optimization. Mean-variance optimization strategy, as part of Modern Portfolio Theory (MPT) 35, constructs portfolios that maximize the expected return for a given level of risk. It uses expected asset returns and covariances to solve an optimization problem. We typically use the past one year’s daily price data to calculate expected returns and covariance matrix. We limit individual stock weights to a maximum of 5%. Mean-variance optimization is a foundational technique in portfolio management and can serve as a classical financial optimization strategy baseline.
5.1.2 Ensemble Trading Agents
Ensemble methods have shown effectiveness in enhancing overall performance by combining selected actions or action probabilities from component RL algorithms 36. We train ensemble trading agents to mitigate policy instability.
@@ 5.1.2 Ensemble Trading Agents
Using KL divergence in objective functions. To enforce diversity among the component agents, we introduce a Kullback-Leibler (KL) divergence term into the agent’s objective function. The KL divergence measures how one probability distribution diverges from another perez2008kldivergence. The KL divergence term penalizes similarities in policies between different agents, encouraging them to converge on different trading strategies. The new objective function for a component agent is as follows:
$\max\penalty 10000\ \penalty 10000\ L_{\text{new}}(\theta_{A})=L_{\text{original}}(\theta_{A})+\lambda\sum_{A\neq B}\text{KL}(\pi_{\theta_{B}}||\pi_{\theta_{A}}),$ (5)
@@
Using various datasets. The financial datasets used for training component agents are varied. For each stock or cryptocurrency, a random percentage change ranging from $1\%$ to $1\%$ is generated and applied to its prices, which shifts the price scale while preserving the original price trends. Agents are also trained on different stocks from the test set for the stock trading task. It enables agents to learn various strategies for a broader range of stocks rather than reacting to a limited number of stocks.
@@
    $\text{Sharpe Ratio}=\frac{\bar{r}_{p}-r_{f}}{\sigma_{p}},$ (6)
@@
Cryptocurrency Trading Task
For cryptocurrency trading at a relatively high frequency, market movements can be modeled as discrete events, which require a discrete action space. As shown in Fig. 5 (b), DQN 37, Double DQN doubledqn, and Dueling DQN 38 are used to handle this discrete action space. In addition, the dataset for a single cryptocurrency is relatively small. DQN and its variants, with fewer parameters and simpler architectures, can be trained faster to avoid overfitting. Moreover, trading at a high frequency requires fast responses, and DQN agents can offer lower latency in decision-making compared to more complex models. The ensemble model uses majority voting to combine the actions of component agents. Majority voting ensures the chosen action reflects consensus among agents, mitigating biases from any single agent’s actions 39. The process is as follows:
@@ 5.2 Evaluation Protocol
  • Backtesting. We use backtesting to evaluate models for trading tasks in FinRL Contests 2024 and 2025, as shown in Fig. 6 (a). The datasets released to participants are used for their training and validation. The withheld out-of-sample dataset is used for evaluation. We did not use the rolling window framework because the evaluation datasets for stock trading have long time spans, and the cryptocurrency data is high-frequency. In both cases, retraining models frequently would be too time-consuming and difficult to manage.
  • Rolling windows. We use the rolling window framework to evaluate models for FinRL Contest 2023 data-centric stock trading and paper trading, as discussed in Sec 4.1. Fig. 6 (b) presents the mechanism of each rolling window for the stock trading task. There are in total $9$ days in a rolling window. We divide it into three stages, namely, $6$ days of training, $2$ days of validation, and $1$ day of trading using the trained DRL agent. The models are retrained for $8$ days before each trading day. The validation stage allows adjusting the hyper-parameters or selecting an agent from the ensemble. Finally, the well-trained agent performs stock trading on the last day of a rolling window.
@@ 6.1 Performance Metrics
We use the well-established quantitative metrics 2 in finance to evaluate models, as shown in Table 3. Specifically, we include cumulative and annualized returns to measure overall profitability and incorporate risk-adjusted performance metrics such as the Sharpe, Sortino, and Calmar ratios. Additionally, we consider downside risk indicators including annualized volatility and maximum drawdown. For each metric, we report its definition, numerical range, and optimization direction (i.e., whether higher or lower values indicate better performance).
Metric Defintiion Range Direction
Cumulative return The total return generated by the trading strategy over a trading period. $(-\infty,+\infty)$ Higher is better
Annualized return The geometric average amount of money earned by the agent each year over a given time period. $(-\infty,+\infty)$ Higher is better
Annualized volatility The annualized standard deviation of daily returns. $0,+\infty)$ Lower is better
Sharpe ratio The excess return per unit of volatility. $(-\infty,+\infty)$ Higher is better
Maximum drawdown The largest single drop in the portfolio value from peak to trough. [1,0][-1,0] Higher is better
Rachev Ratio A risk-adjusted return that measures the potential upside gain compared to potential downside risk, using the tails of the return distribution. $0,+\infty)$ Higher is better
Return over maximum drawdown (RoMaD) It is calculated as the cumulative return divided by the absolute value of maximum drawdown. $(-\infty,+\infty)$ Higher is better
Sortino ratio The excess return divided by the downside deviation. $(-\infty,+\infty)$ Higher is better
Calmar ratio The annualized excess return divided by the maximum drawdown. $(-\infty,+\infty)$ Higher is better
Omega ratio It compares the probability of achieving returns above a threshold to the probability of falling below it. $0,+\infty)$ Higher is better
Win/Loss ratio It is calculated by dividing the number of winning trades by the number of losing trades. $0,+\infty)$ Higher is better
@@ 6.2.1 Data-Centric Stock Trading
The data-centric stock trading task in FinRL Contest 2023 is to train a trading agent by using novel data and feature engineering strategies. The training dataset is the OHLCV dataset with 10 market indicators for all 30 constituent stocks in the Dow Jones Index from 07/01/2010 to 10/24/2023 ($3,352$ trading days). The market indicators are shown in Table 2. The models are evaluated for two periods, 10/25/2023-11/14/2023 (15 trading days, pre submission deadline) and 11/15/2023-11/12-2023 (6 trading days, post submission deadline). Table 4 shows the cumulative return, Sharpe ratio, and maximum drawdown of the top three winners and the Dow Jones Index. The teams used different approaches ******GitHub repo for models and reports: https://github.com/Open-Finance-Lab/FinRL_Contest_2023/tree/main/Task_1:
  • SZU-Fin-621 employed PPO-Switch, which is an ensemble method combining multiple PPO agents. The agent pool was created by training different PPO agents with diverse online portfolio selection (OPS) 40 price features. The agent was selected based on its short-term and long-term returns, and the action space was sparsified to enhance capital efficiency 41.
@@
  • WeCan added 101 formulaic alpha factors 30 as new features in the state, in addition to the provided market indicators.
The results show that the teams’ strategies achieved better risk-adjusted returns and risk management but show poorer profitability compared with the Dow Jones Index. In the first testing period, from 10/25/2023 to 11/14/2023, all teams outperform the Dow Jones Index in Sharpe ratio and maximum drawdown. Nik-Elena achieved the highest Sharpe ratio of 9.56 and the best maximum drawdown of -0.40%, indicating excellent risk-adjusted performance and effective downside protection. In the second testing period from 11/15/2023-11/22/2023, SZU-FIN-621 and WeCan had negative cumulative returns and Sharpe ratios. Nik-Elena had a small positive return of 0.04% but still underperformed the market index. However, Nik-Elena still achieves a higher Sharpe ratio and maximum drawdown than the Dow Jones Index, indicating better risk control even in challenging market environments. Participants’ models showed strong risk management compared to the DJIA in the first testing period, but their generalization to new, unseen market conditions remains a challenge.
@@ 6.2.2 Stock Trading with Ensemble Methods
Ensemble methods. The first method (Ensemble 1) consists of 11 PPO, $1$ SAC, and $1$ DDPG agents; the second method (Ensemble 2) consists of $5$ PPO, $5$ SAC, and $5$ DDPG agents; the third method (Ensemble 3) consists of $10$ agents for each type. As in Section 5.1.2, all three ensemble models use the weighted average approach to combine component agent action probabilities.
Results. As seen in Table 5, the PPO agent achieves the highest cumulative returns of $63.37\%$, Sharpe ratio of $1.55$, and Sortino ratio of $2.44$, showing an ability to maintain high returns with controlled volatility and downside risk. Although DDPG’s cumulative returns are comparable to PPO’s, its higher maximum drawdown of $13.15\%$ signals a greater risk of large value drops, which is a concern for risk management. SAC has a lower maximum drawdown than DDPG but underperforms in other metrics. All individual agents significantly outperform two traditional baselines across all metrics. The ensemble models also maintain profitability and risk management advantages over the baselines. Ensemble 1 has a high cumulative return of $62.60\%$, and as shown in Fig. 7 (a), it shows superior performance from Sep 2022 to Oct 2023. Ensemble 1 also achieves the smallest maximum drawdown and a higher Sharpe ratio than SAC and DDPG. Ensemble 1 and 2 have high RoMaD and Calmar ratios, showing an ability to quickly recover from peak-to-trough losses and a potential for steady growth in market adversities.
Figure 7: Cumulative returns of different strategies for the stock trading task and cryptocurrency trading task.
6.2.3 LLM-generated Signals for Stock Trading
We designed two tasks around LLM-generated signals from news, including FinRL Contest 2024 LLM-Engineered Signals with RLMF and FinRL Contest 2025 FinRL-DeepSeek for Stock Trading.
FinRL Contest 2024 LLM-Engineered Signals with RLMF. Participants fine-tuned LLMs to generate signals through reinforcement learning from market feedback. The training dataset was daily OHLCV data and financial news for $K=7$ large-cap stocks from 01/01/2020 to 10/08/2022 (447 trading days). The effectiveness of LLM-generated signals is evaluated on a dataset from 10/09/2022 to 12/15/2023 (229 trading days), using a long-short trading strategy, i.e., long (buy) the three stocks with the highest sentiment scores and short (sell) the bottom three, closing all positions three days later. Aethernet42††††††https://github.com/Arnav-Gr0ver/ICAIF_FinRL-2024 fine-tuned the LLaMA-3.2-3B-Instruct model through RLMF to generate sentiment scores from news. The team designed a reward function to assess the sentiment score based on market data. The model had a cumulative return of 134.05%, while a simple buy-and-hold strategy yields a 72.71% return. It indicates that the LLM-generated signals were effective even using a simple long-short execution strategy. The win/loss ratio of 1.5 suggests that the model generated more profitable trades than losing ones.
@@ 6.2.3 LLM-generated Signals for Stock Trading
Table 6: The performance of top winner teams for FinRL Contest 2025 Task 1 FinRL-DeepSeek for Stock Trading. The Rachev ratio of Queen’s Gambit are reported wrongly and need further scrutinization.
@@
FinRL Contest 2025 FinRL-DeepSeek for Stock Trading. This task uses a different approach to combine FinRL and LLM-generated signals, where the signals are integrated into the environment to train FinRL agents. The training dataset is daily OHLCV data with 10 market indicators (listed in Table 2) and financial news for Nasdaq 100 constituent stocks from 01/01/2013 to 12/31/2018 (1510 trading days). The models are evaluated on the testing dataset from 01/01/2019 to 12/31/2023 (1258 trading days). The winning teams used different approaches:
@@
  • Otago Alpha‡‡‡‡‡‡https://github.com/lishenjian1996/FINRL_Contest_2025_Task1_Otago_Alpha added the put-call ratio to the feature vector, which is an assessment of put and call option transaction volumes. The put-call ratio was sourced from the OptionMetrics via Wharton Research Data Services (WRDS) database and Bloomberg Terminal.
@@
  • Ruijian & Sally******https://github.com/Ruijian-Zha/FinRL-DAPO-SR used the Group Relative Policy Optimization (GRPO) 43 algorithm, which is refined by Decoupled Clipping and Dynamic sAmpling Policy Optimization (DAPO). The team also designed a reward function adjusted by sentiment and risk.
@@
  • Kuxlnw*†*†*†https://github.com/Vorakorn1001/FinRL-DeepSeek adopted a dynamic model selection mechanism that adapts to market conditions and macroeconomic indicators. The team added new features in the environment, including interest rates, gold prices, and oil prices. They also used DeepSeek-generated sentiment scores and risk levels in the features. They trained specialized FinRL agents for each stock. Then they employed a mechanism to dynamically allocate the most suitable trading agents based on bull or bear markets. The algorithms they use include PPO with 100 training epochs (PPO 100), PPO with DeepSeek-generated signals and 100 training epochs (PPO-DeepSeek 100), CPPO with 100 training epochs (CPPO 100), CPPO with DeepSeek-generated signals and 100 training epochs (CPPO-DeepSeek 100), and Adaptive Portfolio applying the dynamic model selection mechanism. Queen’s Gambit*‡*‡*‡https://github.com/sahar-arshad/QueensGambit-FinRL2025 proposed a two-phase framework, Market-Informed Sentiment for Trading (MIST) to extract signals from news. In the first phase, it used a structured prompt with few-shot examples for DeepSeek-R1 7B to evaluate financial news and assign a sentiment score (1–5). In the second stage, the team designed an advanced prompt incorporating market behavior (i.e., short-term stock price change direction) to either reinforce or contradict the sentiment extracted in the first stage. The final trading signal is integrated in the environment to train FinRL agents. The team trained two PPO agents, one with only the first stage of MIST (PPO with MIST 1) and one with two stages of MIST (PPO with MIST 2).
@@
The performance of the winning teams are shown in Table 6. Team Otago Alpha, Ruijian & Sally, and Queen’s Gambit showed a superior performance over S&P 500 and Nasdaq-100 in cumulative return and Sharpe ratio. Otago Alpha achieved the highest Sharpe ratio ($1.08$) and maximum drawdown ($28/16\%$), and the cumulative return outperformed two market indices, showing a good balance between profitability and risk management. Although Ruijian & Sally and Queen’s Gambit showed high cumulative returns, they had a worse maximum drawdown of $50.24\%$ and $92.47\%$ respectively compared with two market indices, indicating a larger downside risk. A Rachev ratio larger than 1 indicates that their strategies’ upside tail rewards outweighed their downside tail risks during extreme market movements. Kuxlnw’s adaptive portfolio underperformed its PPO 100 model in cumulative return, Sharpe ratio, and Rachev ratio, suggesting that a multi-asset trading agent may have more stable performance in this setting.
6.3 Cryptocurrency Trading Tasks

### v4

@@ 4.2 Market Environments
The financial data is processed into the standard gym-style market environments to build a standard practical environment. This part describes the market environments with near-real market constraints. To address the sampling bottleneck of the training stage, we develop massively parallel market environments on GPUs.
@@ 4.2.1 Market Environment with Trading Constraints
After processing the financial data into standard gym-style market environments, we define the key operations in the FinRL market environment, including:
@@ 4.2.2 Massively Parallel Simulations on GPUs
  • $N$ is the number of parallel SubEnvs.
@@
Improved Sampling Speed with Massively Parallel Environments on GPU. We evaluated the sampling speed measured in samples per second using vectorized environments for stock trading. We used the PPO agent and the OHLCV data of $30$ constituent stocks in the Dow Jones index, from 2020-01-01 to 2022-01-01. The NVIDIA A100 GPU is used. The numbers of parallel environments vary from $1$, $2$, $4$, $\ldots$, to $2,048$. As shown in Fig. 4, the average sampling speed with $2,048$ parallel environments is $227,212.54$ samples per second. The sampling speed is improved by 1,649.93×1,649.93$\times$1 , 649.93 × compared with a single environment. The sampling speed scales approximately linearly with the number of parallel environments. The results show the effectiveness of massively parallel simulation in improving sampling speed in FinRL Contest tasks.
@@ 5.1.1 Market Index and Mean-Variance Method
  • Mean-variance optimization. Mean-variance optimization strategy, as part of Modern Portfolio Theory (MPT) 34, constructs portfolios that maximize the expected return for a given level of risk. It uses expected asset returns and covariances to solve an optimization problem. We typically use the past one year’s daily price data to calculate expected returns and the covariance matrix. We limit individual stock weights to a maximum of 5%. Mean-variance optimization is a foundational technique in portfolio management and can serve as a classical financial optimization strategy baseline.
5.1.2 Ensemble Agent
Ensemble methods have shown effectiveness in enhancing overall performance by combining selected actions or action probabilities from component RL algorithms  35. We train ensemble trading agents to mitigate policy instability.
@@ 5.1.2 Ensemble Agent
Using KL divergence in objective functions. To enforce diversity among the component agents, we introduce a Kullback-Leibler (KL) divergence term into the agent’s objective function. The KL divergence measures how one probability distribution diverges from another  36. The KL divergence term penalizes similarities in policies between different agents, encouraging them to converge on different trading strategies. The new objective function for a component agent is as follows:
$\max\leavevmode\nobreak\ \leavevmode\nobreak\ L_{\text{new}}(\theta_{A})=L_{\text{original}}(\theta_{A})+\lambda\sum_{A\neq B}\text{KL}(\pi_{\theta_{B}}||\pi_{\theta_{A}}),$ (5)
@@
Using various datasets. The financial datasets used for training component agents are varied. For each stock or crypto, a random percentage change ranging from $-1\%$ to $1\%$ is generated and applied to its prices, which shifts the price scale while preserving the original price trends. Agents are also trained on different stocks from the test set for the stock trading task. It enables agents to learn various strategies for a broader range of stocks rather than reacting to a limited number of stocks.
@@
    Sharpe $\text{Sharpe Ratio}=\frac{\bar{r}_{p}-r_{f}}{\sigma_{p}},$ (6)
@@
Crypto Trading Task
For crypto trading at a relatively high frequency, market movements can be modeled as discrete events, which require a discrete action space. As shown in Fig. 5 (b), DQN  37, Double DQN  38, and Dueling DQN  39 are used to handle this discrete action space. In addition, the dataset for a single crypto is relatively small. DQN and its variants, with fewer parameters and simpler architectures, can be trained faster to avoid overfitting. Moreover, trading at a high frequency requires fast responses, and DQN agents can offer lower latency in decision-making compared to more complex models. The ensemble model uses majority voting to combine the actions of component agents. Majority voting ensures the chosen action reflects consensus among agents, mitigating biases from any single agent’s actions  40. The process is as follows:
@@ 5.2 Evaluation Protocol
  • Backtesting. We use backtesting to evaluate models for trading tasks in FinRL Contests 2024 and 2025, as shown in Fig. 6 (a). The datasets released to participants are used for their training and validation. The withheld out-of-sample dataset is used for evaluation. We did not use the rolling window framework because the evaluation datasets for stock trading have long time spans, and the crypto data is high-frequency. In both cases, retraining models frequently would be too time-consuming and difficult to manage.
  • Rolling windows. We use the rolling window framework to evaluate models for FinRL Contest 2023 data-centric stock trading and paper trading, as discussed in Section 4.1. Fig. 6 (b) presents the mechanism of each rolling window for the stock trading task. There are in total $9$ days in a rolling window. We divide it into three stages, namely, $6$ days of training, $2$ days of validation, and $1$ day of trading using the trained DRL agent. The models are retrained for $8$ days before each trading day. The validation stage allows adjusting the hyperparameters or selecting an agent from the ensemble. Finally, the well-trained agent performs stock trading on the last day of a rolling window.
@@ 6.1 Performance Metrics
We use the well-established quantitative metrics 2 in finance to evaluate models, as shown in Table 3. Specifically, we include cumulative and annualized returns to measure overall profitability and incorporate risk-adjusted performance metrics such as the Sharpe, Sortino, and Calmar ratios. Additionally, we consider downside risk indicators, including annualized volatility and maximum drawdown. For each metric, we report its definition, numerical range, and optimization direction (i.e., whether higher or lower values indicate better performance).
Metric Definition Range Direction
Cumulative return The total return generated by the trading strategy over a trading period. (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) Higher is better
Annualized return The geometric average amount of money earned by the agent each year over a given time period. (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) Higher is better
Annualized volatility The annualized standard deviation of daily returns. $[0,+\infty)$ Lower is better
Sharpe ratio The excess return per unit of volatility. (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) Higher is better
Maximum drawdown The largest single drop in the portfolio value from peak to trough. $[-1,0]$ Higher is better
Rachev Ratio A risk-adjusted return that measures the potential upside gain compared to potential downside risk, using the tails of the return distribution. (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) Higher is better
Return over maximum drawdown (RoMaD) It is calculated as the cumulative return divided by the absolute value of maximum drawdown. (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) Higher is better
Sortino ratio The excess return divided by the downside deviation. (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) Higher is better
Calmar ratio The annualized excess return divided by the maximum drawdown. (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) Higher is better
Omega ratio It compares the probability of achieving returns above a threshold to the probability of falling below it. $[0,+\infty)$ Higher is better
Win/Loss ratio It is calculated by dividing the number of winning trades by the number of losing trades. $[0,+\infty)$ Higher is better
@@ 6.2.1 Data-Centric Stock Trading
The data-centric stock trading task in FinRL Contest 2023 is to train a trading agent by using novel data and feature engineering strategies. The training dataset is the OHLCV dataset with 10 market indicators for all 30 constituent stocks in the Dow Jones Index from 07/01/2010 to 10/24/2023 ($3,352$ trading days). The market indicators are shown in Table 2. The models are evaluated for two periods: 10/25/2023-11/14/2023 (15 trading days, pre-submission deadline) and 11/15/2023-11/12-2023 (6 trading days, post-submission deadline). Table 4 shows the cumulative return, Sharpe ratio, and maximum drawdown of the top three winners and the Dow Jones Index. The teams used different approaches ††††††GitHub repo for models and reports: https://github.com/Open-Finance-Lab/FinRL_Contest_2023/tree/main/Task_1:
  • SZU-Fin-621 employed PPO-Switch, which is an ensemble method combining multiple PPO agents. The agent pool was created by training different PPO agents with diverse online portfolio selection (OPS) 41 price features. The agent was selected based on its short-term and long-term returns, and the action space was sparsified to enhance capital efficiency 42.
@@
  • WeCan added 101 formulaic alpha factors 29 as new features in the state, in addition to the provided market indicators.
The results show that the teams’ strategies achieved better risk-adjusted returns and risk management, but showed poorer profitability compared with the Dow Jones Index. In the first testing period, from 10/25/2023 to 11/14/2023, all teams outperformed the Dow Jones Index in the Sharpe ratio and maximum drawdown. Nik-Elena achieved the highest Sharpe ratio of 9.56 and the best maximum drawdown of -0.40%, indicating excellent risk-adjusted performance and effective downside protection. In the second testing period from 11/15/2023 to 11/22/2023, SZU-FIN-621 and WeCan had negative cumulative returns and Sharpe ratios. Nik-Elena had a small positive return of 0.04% but still underperformed the market index. However, Nik-Elena still achieves a higher Sharpe ratio and maximum drawdown than the Dow Jones Index, indicating better risk control even in challenging market environments. Participants’ models showed strong risk management compared to the DJIA in the first testing period, but their generalization to new, unseen market conditions remains a challenge.
@@ 6.2.2 Stock Trading with Ensemble Methods
Ensemble methods. The first method (Ensemble 1) consists of $1$ PPO, $1$ SAC, and $1$ DDPG agents; the second method (Ensemble 2) consists of $5$ PPO, $5$ SAC, and $5$ DDPG agents; and the third method (Ensemble 3) consists of $10$ agents for each type. As in Section 5.1.2, all three ensemble models use the weighted average approach to combine component agent action probabilities.
Results. As seen in Table 5, the PPO agent achieves the highest cumulative returns of $63.37\%$, Sharpe ratio of $1.55$, and Sortino ratio of $2.44$, showing an ability to maintain high returns with controlled volatility and downside risk. Although DDPG’s cumulative returns are comparable to PPO’s, its higher maximum drawdown of $-13.15\%$ signals a greater risk of large value drops, which is a concern for risk management. SAC has a lower maximum drawdown than DDPG but underperforms in other metrics. All individual agents significantly outperform two traditional baselines across all metrics. The ensemble models also maintain profitability and risk management advantages over the baselines. Ensemble 1 has a high cumulative return of $62.60\%$, and as shown in Fig. 7 (a), it shows superior performance from Sep 2022 to Oct 2023. Ensemble 1 also achieves the smallest maximum drawdown and a higher Sharpe ratio than SAC and DDPG. Ensembles 1 and 2 have high RoMaD and Calmar ratios, showing an ability to quickly recover from peak-to-trough losses and a potential for steady growth in market adversities.
Figure 7: Cumulative returns of different strategies for the stock trading task and crypto trading task.
6.2.3 LLM-engineered Signals for Stock Trading
We designed two tasks around LLM-engineered signals from news, including FinRL Contest 2024 LLM-Engineered Signals with RLMF and FinRL Contest 2025 FinRL-DeepSeek for Stock Trading.
FinRL Contest 2024 LLM-Engineered Signals with RLMF. Participants fine-tuned LLMs to generate signals through reinforcement learning from market feedback. The training dataset was daily OHLCV data and financial news for $K=7$ large-cap stocks from 01/01/2020 to 10/08/2022 (447 trading days). The effectiveness of LLM-engineered signals is evaluated on a dataset from 10/09/2022 to 12/15/2023 (229 trading days), using a long-short trading strategy, i.e., long (buy) the three stocks with the highest sentiment scores and short (sell) the bottom three, closing all positions three days later. Aethernet‡‡‡‡‡‡https://github.com/Arnav-Gr0ver/ICAIF_FinRL-202443 fine-tuned the LLaMA-3.2-3B-Instruct model through RLMF to generate sentiment scores from news. The team designed a reward function to assess the sentiment score based on market data. The model had a cumulative return of 134.05%, while a simple buy-and-hold strategy yields a 72.71% return. It indicates that the LLM-engineered signals were effective even using a simple long-short execution strategy. The win/loss ratio of 1.5 suggests that the model generated more profitable trades than losing ones.
@@ 6.2.3 LLM-engineered Signals for Stock Trading
Table 6: The performance of top winner teams for FinRL Contest 2025 Task 1 FinRL-DeepSeek for Stock Trading. The Rachev ratio of Queen’s Gambit is reported wrongly and needs further scrutiny.
@@
FinRL Contest 2025 FinRL-DeepSeek for Stock Trading. This task uses a different approach to combine FinRL and LLM-engineered signals, where the signals are integrated into the environment to train FinRL agents. The training dataset is daily OHLCV data with 10 market indicators (listed in Table 2) and financial news for Nasdaq 100 constituent stocks from 01/01/2013 to 12/31/2018 (1510 trading days). The models are evaluated on the testing dataset from 01/01/2019 to 12/31/2023 (1258 trading days). The winning teams used different approaches:
@@
  • Otago Alpha 44 added the put-call ratio to the feature vector, which is an assessment of put and call option transaction volumes. The put-call ratio was sourced from the OptionMetrics via the Wharton Research Data Services (WRDS) database and Bloomberg Terminal.
@@
  • Ruijian & Sally******https://github.com/Ruijian-Zha/FinRL-DAPO-SR45 used the Group Relative Policy Optimization (GRPO) 46 algorithm, which is refined by Decoupled Clipping and Dynamic sAmpling Policy Optimization (DAPO). The team also designed a reward function adjusted by sentiment and risk.
@@
  • Kuxlnw*†*†*†https://github.com/Vorakorn1001/FinRL-DeepSeek47 adopted a dynamic model selection mechanism that adapts to market conditions and macroeconomic indicators. The team added new features to the environment, including interest rates, gold prices, and oil prices. They also used DeepSeek-generated sentiment scores and risk levels in the features. They trained specialized FinRL agents for each stock. Then they employed a mechanism to dynamically allocate the most suitable trading agents based on bull or bear markets. The algorithms they use include PPO with 100 training epochs (PPO 100), PPO with DeepSeek-generated signals and 100 training epochs (PPO-DeepSeek 100), CPPO with 100 training epochs (CPPO 100), CPPO with DeepSeek-generated signals and 100 training epochs (CPPO-DeepSeek 100), and Adaptive Portfolio applying the dynamic model selection mechanism.
  • Queen’s Gambit*‡*‡*‡https://github.com/sahar-arshad/QueensGambit-FinRL202548 proposed a two-phase framework, Market-Informed Sentiment for Trading (MIST), to extract signals from news. In the first phase, it used a structured prompt with few-shot examples for DeepSeek-R1 7B to evaluate financial news and assign a sentiment score (1–5). In the second stage, the team designed an advanced prompt incorporating market behavior (i.e., short-term stock price change direction) to either reinforce or contradict the sentiment extracted in the first stage. The final trading signal is integrated into the environment to train FinRL agents. The team trained two PPO agents, one with only the first stage of MIST (PPO with MIST 1) and one with two stages of MIST (PPO with MIST 2).
@@
The performance of the winning teams are shown in Table 6. Team Otago Alpha, Ruijian & Sally, and Queen’s Gambit showed a superior performance over S&P 500 and Nasdaq-100 in cumulative return and Sharpe ratio. Otago Alpha achieved the highest Sharpe ratio ($1.08$) and maximum drawdown ($-28/16\%$), and the cumulative return outperformed two market indices, showing a good balance between profitability and risk management. Although Ruijian & Sally and Queen’s Gambit showed high cumulative returns, they had a worse maximum drawdown of $-50.24\%$ and $-92.47\%$ respectively compared with two market indices, indicating a larger downside risk. A Rachev ratio larger than 1 indicates that their strategies’ upside tail rewards outweighed their downside tail risks during extreme market movements. Kuxlnw’s adaptive portfolio underperformed its PPO 100 model in cumulative return, Sharpe ratio, and Rachev ratio, suggesting that a multi-asset trading agent may have more stable performance in this setting.
6.3 Crypto Trading Tasks